THE ^{14}N/^{15}N ISOTOPE RATIO IN DENSE MOLECULAR CLOUDS

\underline{G. ADANDE}, L.M. ZIURYS, Department of Chemistry, Steward Observatory, University of Arizona, Tucson.

Observations of the 3-2 and 1-0 transitions of HNC at 271.981 and 90.663 GHz, H^{15}NC at 266.567 and 88.865 GHz were conducted using the Arizona Radio Observatory (ARO) 12m at 2 and 3mm in wavelength and Submillimeter Telescopes (SMT) at 1mm. The ^{13}C isotopomer was also observed to evaluate the high optical depth in the main isotope, hence the 3-2, 2-1, 1-0 transitions of HN^{13}C at 261.263, 174.179 and 87.090 GHz were recorded as well. Observations were made toward the molecular clouds SgrB2, W31, G34.3, W51M, M17-SW, DR-21, L134, Orion A, W3 (OH), NGC7538 and S156, located at various distances from the Galactic Center. The preliminary results indicate a ratio ^{14}N/^{15}N of 120-200, similarly to the values found in comets.